OpenStudy (anonymous):

Half Angle Identity: tan7/12

OpenStudy (anonymous):

@calculusfunctions No worries man, if you have to leave, by all means don't worry about me. I'll find help on this material.

OpenStudy (calculusfunctions):

Do you mean\[\tan \frac{ 7\pi }{ 12 }\]

OpenStudy (anonymous):

It doesn't show that in my notes but that would make more sense

OpenStudy (calculusfunctions):

Half-angle identity:\[\tan \frac{ \theta }{ 2 }=\sqrt{\frac{ 1-\cos \theta }{ 1+\cos \theta }}\]

OpenStudy (anonymous):

tan(105) = \[\tan(\alpha/2)\]

OpenStudy (anonymous):

so \[\alpha = 210\]

OpenStudy (calculusfunctions):

Yes\[\frac{ 7\pi }{ 12 }=105° \]so that θ or α = 210° as you said. Very good!

OpenStudy (anonymous):

\[\tan(\alpha/2)=\sqrt{\frac{ 1-\cos(210) }{ 1+\cos(210) }}\]

OpenStudy (calculusfunctions):

Excellent! Now simplify.

OpenStudy (anonymous):

\[-2-\sqrt{3}\] ??

OpenStudy (anonymous):

Sorry kinda guessed at that

OpenStudy (calculusfunctions):

Nope, Try again. Remember that cos (210°) = -(√3)/2

OpenStudy (anonymous):

\[2+\sqrt{3}\]

OpenStudy (calculusfunctions):

No, do you want me to type out the steps so that you can study them later?

OpenStudy (anonymous):

Yes, please, I'm not seeing this part

OpenStudy (calculusfunctions):

\[-\sqrt{\frac{ 1-\cos (210°) }{ 1+\cos (210°) }} \]\[=-\sqrt{\frac{ 1-(\frac{ -\sqrt{3} }{ 2 }) }{ 1+(\frac{ -\sqrt{3} }{ 2 }) }}\]\[=-\sqrt{\frac{ 1+\frac{ \sqrt{3} }{ 2 } }{ 1-\frac{ \sqrt{3} }{ 2 } }}\]\[=-\sqrt{\frac{ \frac{ 2+\sqrt{3} }{ 2 } }{ \frac{ 2-\sqrt{3} }{ 2 } }}\]\[=-\sqrt{\frac{ 2+\sqrt{3} }{ 2-\sqrt{3} }}\]\[=\sqrt{\frac{ (2+\sqrt{3}) }{ (2-\sqrt{3}) }⋅\frac{ (2+\sqrt{3}) }{ (2+\sqrt{3}) }}\]\[=-\sqrt{\frac{ 4+4\sqrt{3}+3 }{ 4-3 }} \]\[=-\sqrt{\frac{ 7+4\sqrt{3} }{ 1 }}\]\[=-\sqrt{7+4\sqrt{3}} \]

OpenStudy (calculusfunctions):

The reason for the negative is because 105° is in the second quadrant where tangent is negative.

OpenStudy (calculusfunctions):

I hope that helps. I have to go so catch you later.

OpenStudy (anonymous):

Thanks a ton for all the help, it makes a lot more sense now. Later.

OpenStudy (calculusfunctions):

Later. Glad to know you understand.